(a) Find the change in entropy when 2 kg of Oxygen at 600C are (30) mixed...

  1. (a) Find the change in entropy when 2 kg of Oxygen at 600C are (30) mixed with 6 kg of Nitrogen at 400C in a rigid vessel. The
    initial pressure of each component is 103 kPa and remains
    the same for that of the mixture.

    (b) What is the amount of heat transferred during the mixing process?

    Given, Cv = 0.743 J/g.oK for Nitrogen and Cv = 0.659 J/g.oK for Oxygen.
Answer #1

२ Jl. K Cue 0.743 I + 32 9 mol Steb: 1 Cuiven, Mass of anygen, mar 2 kg Heat capacity at constant volume, G64433 M? Holes ofR = Universal Gas Constant 1 p = G+ 20 G = 23.7 + 8.314 Go . 39.01 J/moln K oor temperature of O2, To, 600°C boot 273 Ton 873G. 0.659 Inglik + 28 g/mol G. 18.45 Il malik G - G+R 18.45 +.8.314 96.74 Il mol. K 2 temperature of Na TNG 2 400 °C 2 400+ 27Step Let t be the Equilibrium temperature of the mixture by Taking Energy Balance, Ton - T) (T-Tren] M107 MN CPN 2 cror 62.5For an Ideal Gas, Enkopy change of System to as) System & lm Tz - R lm Pz Т) en P at constant pressure G = Heat Capacity T2,nor Og Com Ton 62.5 mole * 32:01 Il mol: K 2 214.28 male & 26.74 /mol K Here Pope [constant Messure process] lim P₂ zo Enkopy(As) 424.87 YIK Na Total Entropy change of mining (15) - (As)o + (AS) 2 -372.2 + 424.87 52.67 Ilk 2 (b) Heat transferred Equa